Sentinel Company is considering an investment in technology to improve its operations. The investment will require an initial outlay of $257,000 and will yield the following expected cash flows. Management requires investments to have a payback period of 3 years, and it requires a 8% return on investments. (PV of $1, FV of $1, PVA of $1, and FVA of $1) (Use appropriate factor(s) from the table provided.)

Period Cash Flow
1 $ 48,700
2 52,100
3 76,600
4 95,600
5 125,800


Required:
1. Determine the payback period for this investment.
2. Determine the break-even time for this investment.
3. Determine the net present value for this investment.
Answer #1

Correct Answer:

Requirement 1: Payback period: = 3.83 years.

Working:

Payback period = year before full recovery + (unrecovered amount at the start to the period / cash flow during the next period.

Year

Net Cash Flow

Cumulative Cash Flow

0

$ (2,57,000.00)

$ (2,57,000.00)

1

$      48,700.00

$ (2,08,300.00)

2

$      52,100.00

$ (1,56,200.00)

3

$      76,600.00

$     (79,600.00)

4

$      95,600.00

$      16,000.00

5

$   1,25,800.00

$   1,41,800.00

Year before full recovery = 3 years

Unrecovered amount at the start of the next period = $ 79,600

Cash flow during the next period = $ 95,600

Payback period = 3 + (79600/95600)

= 3 + 0.832636

= 3.83 years

Requirement 2: break-even time for the investment (Discounted payback period) = 4.42 years

Break-even time for the investment (Discounted payback period) = year before full recovery + (unrecovered discounted amount at the start to the period / discounted cash flow during the next period.

Year

Net Cash Flow

discount rate @ 8%

Discounted cash flow

Cumulative Cash Flow

0

$ (2,57,000.00)

1

$     (2,57,000.00)

$   (2,57,000.00)

1

$      48,700.00

0.925925926

              45,092.59

$   (2,11,907.41)

2

$      52,100.00

0.85733882

              44,667.35

$   (1,67,240.05)

3

$      76,600.00

0.793832241

              60,807.55

$   (1,06,432.51)

4

$      95,600.00

0.735029853

              70,268.85

$      (36,163.65)

5

$   1,25,800.00

0.680583197

              85,617.37

$       49,453.71

Break-even time for the investment = 4 +( 36163.65/85617.37)

= 4 + 0.42238687

= 4.42 years

Requirement 3: Net present value: $ 49,453.71

Net present value = total present value of all future cash inflows – cash outflow

Year

Net Cash Flow

discount rate @ 8%

Present value fo cash flow (outflow)

0

$ (2,57,000.00)

1

$     (2,57,000.00)

1

$      48,700.00

0.925925926

$           45,092.59

2

$      52,100.00

0.85733882

$           44,667.35

3

$      76,600.00

0.793832241

$           60,807.55

4

$      95,600.00

0.735029853

$           70,268.85

5

$   1,25,800.00

0.680583197

$           85,617.37

Net Present value

$           49,453.71

End of answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks.